# Magnetic reversals in a geodynamo model with a stably-stratified layer

### v1

@@ 4 Results
Stable dipolar dynamos are typically characterised by a dipolar strength $f_{\mathrm{dip}}>0.5$, whereas lower values correspond to multipolar solutions, in which the dipole component is no longer dominant. The transition between these two dynamo solutions remains a major unresolved issue in our understanding of the geodynamo. Figure 3 (a) shows the dipolar-multipolar transition as a function of the Rayleigh number. In the case of a fully convective system ($\mathcal{H}_{s}=0$), the transition takes place at around $\mathrm{Ra}/\mathrm{Ra}_{c}=15$. The introduction of the SSL shifts this transition to larger values, around $\mathrm{Ra}/\mathrm{Ra}_{c}=27.5$ for $\mathcal{H}_{s}=0.24L$, and makes the dipolar-multipolar transition more abrupt. Figure 3 (b) shows the same transition as a function of the local Rossby number defined in Eq. (14). The critical value of $\mathrm{Ro}_{\ell}$ for the transition depends on the size of the layer.
@@
Figure 4: (a) Evolution of the magnetic field in the D-Q phase space for different amplitudes of the heterogeneous heat flux $\delta q$ for the simplest pattern $Y_{1}^{0}$. The points correspond to a portion of the time series ($50\tau_{\nu}$) taken from the statistically steady state of the dynamo solution. (b) Time averaged hemisphericity, defined as in Eq. (15), as a function of $\delta q$. All simulations correspond to $\mathrm{E}=10^{-3}$, $\Pr=1$, and $\mathrm{Pm}=10$. Error bars indicate the standard deviation.
@@
Having established the influence of the SSL on the dipolar-multipolar transition, we now turn to exploring the effects of an imposed axisymmetric heterogeneous heat flux at the CMB. We choose the case with a large SSL with $\mathcal{H}_{s}=0.24L$ and a value of $\mathrm{Ra}/\mathrm{Ra}_{c}=25$ such that it is close to the transition with a stable dipolar solution. We start by imposing the $Y_{1}^{0}$ pattern at the CMB, and vary the amplitude of the symmetry breaking $\delta q$. We verified that imposing a heat flux at the CMB lower than $10\%$, the equatorial symmetry of the flow does not change significantly. Figure 4 shows the phase diagram of the axial dipole $D$ and quadrupole $Q$ components of the magnetic field at the CMB, defined in Eq. (7), for different values of $\delta q$ in the presence of an SSL. In the case of no symmetry breaking (black curve), the magnetic field fluctuates around a stable state with a strong dipole component, and a quadrupole that fluctuates around zero. As the amplitude of the heterogeneous heat flux increases (red curve), the quadrupole is enhanced, and the magnetic field fluctuates around a mixed state, and does not display reversals. Moreover, when the amplitude of the heterogeneous heat flux is negative (blue curve), the quadrupole also changes its sign. This enhancement of the quadrupole component is associated with an increase in the hemisphericity of the magnetic field, defined in Eq. (15) and shown in Fig. 4-(b). It is remarkable that breaking the heat flux symmetry by only a few percent - leading to a minimal modification of the velocity field ($E_{\mathrm{kin}}^{\mathrm{asym}}/E_{\mathrm{kin}}^{\mathrm{tot}}<20$%) - can result in a strongly hemispherical magnetic field.
@@
Figure 5: (a) Evolution of the magnetic field in the D-Q phase space for different amplitudes of the heterogeneous heat flux $\delta q$ for the pattern $Y_{3}^{0}$. Arrows indicate the orientation of the typical trajectories of reversals. (b) Time series of $D$ and $Q$ for the case $\delta q=5.3\%$. All simulations correspond to $\mathrm{E}=10^{-3}$, $\Pr=1$, and $\mathrm{Pm}=10$.
@@
We also investigate the effects of a more complex heterogeneous heat flux pattern at the CMB given by the spherical harmonic $Y_{3}^{0}$. From a geophysical perspective, this choice is motivated by seismic and mantle convection studies suggesting that large-scale, non-uniform heat flux anomalies exist at the CMB driven by localised and asymmetric features. Therefore, exploring the role of a $Y_{3}^{0}$ pattern offers insight into how more realistic CMB heat flux distributions might influence the geodynamo. The results of the simulations are shown in Fig. 5. As the amplitude of the heterogeneous heat flux increases, the system eventually reaches a regime in which there are magnetic reversals, switching from a $\bm{B}\rightarrow-\bm{B}$ solution, where both the dipole and quadrupole components change sign (red curve). Arrows indicate the typical trajectories of reversals. The time series of the dipole and quadrupole for $\delta q=5.3\%$ is shown in Fig. 5 (b). It is found systematically that during the reversal process, the dipole reverses first, and the quadrupole follows slightly later. Reversals follow the opposite trajectory when the amplitude of the heterogeneous heat flux is negative (blue curve). A long evolution of the dipolar angle and the dipolar strength $f_{\mathrm{dip}}$ for $\delta q=5.3\%$ is shown in Fig. 6. In the period of $120$ dipole diffusion times, we observe several non-periodic chaotic reversals of the magnetic field. During the reversal, the dipolar strength decreases from $f_{\mathrm{dip}}\approx 0.9$ to a value $f_{\mathrm{dip}}<0.5$, consistent with the reduction of the dipole component during this event shown in Fig. 5(b). The high value of $f_{\mathrm{dip}}$ between reversals can be attributed to the presence of the SSL, allowing for a strong dipolar magnetic field that reverses, in contrast with the magnetic reversals observed in the absence of a layer in the multipolar regime, dominated by high order modes.
@@
We can also consider a linear combination of the two patterns $Y_{1}^{0}$ and $Y_{3}^{0}$ as a heterogeneous heat flux at the CMB. Figure 7 shows the evolution of the magnetic field in the D-Q phase space for a linear combination of these two patterns, both with the same amplitude and a total flux variation $\delta q=1.75\%$. This heterogeneous heat flux leads to polarity reversals of the magnetic field, with a trajectory that is qualitatively similar to the one described by $Y_{3}^{0}$, but here the time delay between the reversal of the dipole and quadrupole is larger, and the trajectories deviate significantly from the origin. The arrows in the figure show the direction of the reversal. The phenomenology observed here differs from the one described in the absence of a SSL, where both the dipole and quadrupole components fluctuate around the origin, as shown in the inset of Fig. 7(a).
@@ 5 Discussion
Figure 7: Evolution of the dipole and quadrupole magnetic field for a linear combination of the heterogeneous heat flux patterns $Y_{1}^{0}+Y_{3}^{0}$ with an amplitude $\delta q=1.75\%$. The arrows indicate the orientation of reversals. The inset shows a typical simulation with reversals in a fully convective domain. All simulations correspond to $\mathrm{E}=10^{-3}$, $\Pr=1$, and $\mathrm{Pm}=10$.
The phenomenology presented in the previous section, involving the emergence of hemispheric dynamos and polarity reversals due to an equatorial symmetry breaking, can be interpreted using low-dimensional models. These models consider that the axial dipole $D$ and axial quadrupole $Q$ are the two dominant components of the magnetic field, which interact in a non-linear manner coupled by the hydrodynamic properties of the system. Taking into account the $\bm{B}\rightarrow-\bm{B}$ invariance of the MHD equations, this model is written as (Pétrélis and Fauve, 2008; Pétrélis et al., 2009)
@@
where all the parameters $C_{ij}$, $\alpha$, $\beta$ and $\sigma_{D,Q}$ are real numbers. In particular, the coefficients $\sigma_{D}$ and $\sigma_{Q}$ determine the dipole and quadrupole growth rates in the limit in which these two modes are decoupled and the non-linear interactions are negligible. When $\sigma_{D}$ and $\sigma_{Q}$ are similar, an equatorial symmetry breaking will result either in periodic polarity reversals or in hemispheric dynamos (Gallet and Pétrélis, 2009). In the full model, stochastic noise is introduced to account for the effects of turbulence (Pétrélis and Fauve, 2008). Alternatively, Eqs (16) can be coupled to a third mode modelling the asymmetric velocity mode generated by the heterogeneous heat flux, resulting in a deterministic model for chaotic inversions (Gissinger, 2012).
@@
It has been shown that such low-dimensional models based on dipole-quadrupole interactions can reproduce fairly well the magnetic field reversals observed on Earth (Pétrélis et al., 2009, 2011) and in experimental dynamos (Berhanu et al., 2007; Gissinger, 2010). However, this behaviour has never been observed, so far, in numerical modelling of the geodynamo, even in the presence of a heterogeneous heat flux (Gissinger et al., 2012). Interestingly, it is therefore the presence of an SSL in the present simulations that gives rise to this long-anticipated phenomenology. This could be a consequence of the attenuation of fast modes of the magnetic field due to the skin effect in the SSL. A more appealing interpretation involves the relative growth rates of the dipole and quadrupole components and their non-linear interaction. To understand the behaviour of these quantities, we perform DNS of the kinematic dynamo model, which corresponds to the same system of equations (1)-(4) eliminating the Lorentz force from Eq. (1). We first focus on the hydrodynamic properties of the system, in particular on the equatorial symmetry of the flow as a function of the Rayleigh number. Figure 8(a) shows the relative antisymmetric kinetic energy for different values of the size of the SSL. For low values of $\mathrm{Ra}/\mathrm{Ra}_{c}$, this quantity is zero indicating that the flow is symmetric with respect to the equator. In this regime, the dipole and quadrupole components are independent as they are decoupled. At around $\mathrm{Ra}=10\mathrm{Ra}_{c}$, the antisymmetric kinetic energy undergoes a transition and the equatorial symmetry of the flow is broken. When this happens, $D$ and $Q$ are coupled, and the growth rate of these two components are the same, favouring for magnetic reversals. It is interesting to note that this transition is consistent with the dipolar-multipolar transition in the fully convective system, and explains why magnetic reversals are obtained in this regime.
@@
More interestingly, we now focus on the effect of the SSL on the kinematic growth rates of the dipole and quadrupole modes below the symmetry breaking threshold, in the regime where the two linear growth rates remain decoupled. Figure 8(b) shows the dipole and quadrupole growth rates for $\mathrm{Ra}=7\mathrm{Ra}_{c}$ for different values of $\mathcal{H}_{s}$. When the size of the SSL increases, the relative difference between the kinematic growth rates of the dipolar and quadrupolar modes decreases. This behaviour can be understood with the prediction by Proctor (1977), who showed that in $\alpha^{2}$-dynamos surrounded by a perfectly conducting, motionless outer layer, a degeneracy arises between the growth rates of the dipole and quadrupole components. In this context, the SSL beneath the core-mantle boundary can be interpreted as an effective conducting boundary, leading to a similar degeneracy as its thickness increases (Favier and Proctor, 2013). These conditions strongly favour the emergence of the low-dimensional dynamics described above: as the dipole and quadrupole growth rates $\sigma_{D}$ and $\sigma_{Q}$ approach one another, even a weak symmetry breaking induces a spectacularly strong coupling between the two modes, enhancing the dynamics predicted by low-dimensional models - including both chaotic reversals and the development of hemispherical magnetic fields.
@@ 6 Conclusions
In this work, we studied the effects of a stably-stratified layer (SSL) below the core-mantle boundary (CMB) on the morphology of the magnetic field, with particular focus on polarity reversals. Using direct numerical simulations of the incompressible magnetohydrodynamics (MHD) equations in a rotating spherical shell driven by convection, we showed that the SSL enhances the dipolar strength of the magnetic field at the CMB due to the skin effect, that attenuates rapidly fluctuating fields. As a result, the transition from dipolar to multipolar dynamos occurs at higher values of the Rayleigh number when the SSL is present, and the transition becomes more abrupt.
@@
By imposing both an SSL and a heterogeneous heat flux at the CMB to break the equatorial symmetry of the flow, we observed a variety of dynamo regimes, including hemispheric dynamos and polarity reversals. Hemispheric dynamos arise from an increase in the quadrupolar component of the magnetic field, which is driven by the amplitude of the imposed heterogeneous heat flux and results in a weak equatorial symmetry breaking of the flow. For the parameters studied in this work, we found that the simplest heat flux pattern, described by the spherical harmonic $Y_{1}^{0}$, controls the hemisphericity of the magnetic field. In contrast, a more complex pattern $Y_{3}^{0}$ is responsible for triggering polarity reversals. When a linear combination of these two patterns is imposed, reversals follow a clear trajectory in the dipole-quadrupole (DQD-Q) phase space, with the dipole reversing systematically before the quadrupole. This reversal behaviour differs from that observed in fully convective systems within the multipolar regime, where both the dipole and quadrupole components fluctuate without following a well-defined trajectory in the DQD-Q plane.
@@
We interpret these results and the role of the SSL by means of kinematic dynamo simulations. In fully convective dynamos, we associate the dipolar-multipolar transition with a hydrodynamical instability - controlled by the Rayleigh number - that breaks the equatorial symmetry of the flow. When the flow remains symmetric with respect to the equatorial plane, the dipole and the quadrupole become decoupled, and the growth rates of the dipole $\sigma_{D}$ and the quadrupole $\sigma_{Q}$ approach each other as the thickness of the SSL increases. This behaviour is consistent with the prediction by Proctor (1977), who showed that in $\alpha^{2}$-dynamos surrounded by a perfect conductor, a degeneracy arises between the dipolar and quadrupolar growth rates.
@@
This interpretation of the SSL, and its critical role in reducing the difference between dipolar and quadrupolar growth rates ($\sigma_{D}-\sigma_{Q}$), provides an elegant resolution to a long-standing question regarding the robustness of magnetic field reversals. Indeed, despite considerable variability in core conditions - such as inner-core growth, fluctuations in the heat flux, and other long-term evolutionary processes - the presence of a substantial SSL acts as a stabilizing factor. By maintaining the near-degeneracy of $\sigma_{D}$ and $\sigma_{Q}$, it ensures that the low-dimensional dynamics driven by the dipole-quadrupole interaction remain robust, even in the presence of other complex small scale modes. Consequently, even small symmetry breaking effects become sufficient to sustain the observed reversal behaviour over geological timescales, making Earth’s dynamo remarkably resilient to changes in its internal properties. The results shown in this work were obtained for a relatively high value of the Ekman number of $\mathrm{E}=10^{-3}$, which is quite far from realistic values for the Earth of $\mathrm{E}=10^{-15}$. For a future study, we will consider how is the dependence of this phenomenology with the Ekman number and also the magnetic Prandtl number.

### v3

@@ 4.1 Parametric study
Stable dipolar dynamos are typically characterised by a dipolar strength $f_{\mathrm{dip}}>0.5$, whereas lower values correspond to multipolar solutions, in which the dipole component is no longer dominant. The transition between these two dynamo solutions remains a major unresolved issue in our understanding of the geodynamo. Figure 3 (a) shows the dipolar-multipolar transition as a function of the Rayleigh number. In the case of a fully convective system ($\mathcal{H}_{s}=0$), the transition takes place at around $\mathrm{Ra}/\mathrm{Ra}_{c}=15$. The introduction of the SSL shifts this transition to larger values, around $\mathrm{Ra}/\mathrm{Ra}_{c}=27.5$ for $\mathcal{H}_{s}=0.24L$, and makes the dipolar-multipolar transition more abrupt. Figure 3 (b) shows the same transition as a function of the local Rossby number defined in Eq. (15). The critical value of $\mathrm{Ro}_{\ell}$ for the transition depends on the size of the layer. Zaire2022 proposed that the transition between stable and reversing dynamos is controlled by the ratio between kinetic and magnetic energies, that we report in Table 1. Their study in fully convective systems suggests that the transition takes place at around $E_{\mathrm{kin}}/E_{\mathrm{mag}}\approx 0.7$. In our case, as the size of the stably-stratified layer increases, this transition seems to shift to larger values, being around $2$ for $\mathcal{H}_{s}=0.24L$.
@@
Figure 4: (a) Evolution of the magnetic field at the CMB in the D-Q phase space for different amplitudes of the heterogeneous heat flux $\delta q$ for the simplest pattern $Y_{1}^{0}$. The points correspond to a portion of the time series ($50\tau_{\nu}$) taken from the statistically steady state of the dynamo solution. (b) Time averaged hemisphericity, defined as in Eq. (16), as a function of $\delta q$. All simulations correspond to $\mathrm{E}=10^{-3}$, $\Pr=1$, $\mathrm{Pm}=10$, and $\mathrm{Ra}/\mathrm{Ra}_{c}=25$. Error bars indicate the standard deviation.
4.2 Effects of an heterogeneous heat flux at the CMB
@@
Having established the influence of the SSL on the dipolar-multipolar transition, we now turn to exploring the effects of an imposed axisymmetric heterogeneous heat flux at the CMB. We choose the case with a large SSL with $\mathcal{H}_{s}=0.24L$ and a value of $\mathrm{Ra}/\mathrm{Ra}_{c}=25$ such that it is close to the transition with a stable dipolar solution. The input and output quantities of these simulations are summarised in Table 2. We start by imposing the $Y_{1}^{0}$ pattern at the CMB, and vary the amplitude of the symmetry breaking $\delta q$. We verified that when imposing a heat flux at the CMB lower than $10\%$, the equatorial symmetry of the flow does not change significantly. Figure 4 shows the phase diagram of the axial dipole $D$ and quadrupole $Q$ components of the magnetic field at the CMB, defined in Eq. (7), for different values of $\delta q$ in the presence of an SSL. In the case of no symmetry breaking (black curve), the magnetic field fluctuates around a stable state with a strong dipole component, and a quadrupole that fluctuates around zero. As the amplitude of the heterogeneous heat flux increases (red curve), the quadrupole is enhanced, and the magnetic field fluctuates around a mixed state, and does not display reversals. A movie of the simulation with $\delta q=8.5\%$ can be found in the Supplementary Material supp. Moreover, when the amplitude of the heterogeneous heat flux is negative (blue curve), the quadrupole also changes its sign. The straight line connecting to the origin is to indicate the angle between the dipole and quadrupole components. This enhancement of the quadrupole component is associated with an increase in the hemisphericity of the magnetic field, defined in Eq. (16) and shown in Fig. 4-(b). It is remarkable that breaking the heat flux symmetry by only a few percent - leading to a minimal modification of the velocity field ($E_{\mathrm{kin}}^{\mathrm{asym}}/E_{\mathrm{kin}}^{\mathrm{tot}}<20$%) - can result in a strongly hemispherical magnetic field, in line with the results of an $\alpha^{2}$ model of dynamo (Gallet2009).
@@
Figure 5: (a) Evolution of the magnetic field at the CMB in the D-Q phase space for different amplitudes of the heterogeneous heat flux $\delta q$ for the pattern $Y_{3}^{0}$. Arrows indicate the orientation of the typical trajectories of reversals. (b) Time series of $D$ and $Q$ for the case $\delta q=5.3\%$. All simulations correspond to $\mathrm{E}=10^{-3}$, $\Pr=1$, $\mathrm{Pm}=10$, and $\mathrm{Ra}/\mathrm{Ra}_{c}=25$.
@@
We also investigate the effects of a more complex heterogeneous heat flux pattern at the CMB given by the spherical harmonic $Y_{3}^{0}$. This choice is motivated primarily by theoretical and methodological considerations: $Y_{3}^{0}$ provides a simple axisymmetric pattern with increased latitudinal complexity compared to $Y_{1}^{0}$, allowing us to investigate the effects of higher-degree axisymmetric heterogeneity. The results of the simulations are shown in Fig. 5. As the amplitude of the heterogeneous heat flux increases, the system eventually reaches a regime in which there are magnetic reversals, switching from a $\bm{B}\rightarrow-\bm{B}$ solution, where both the dipole and quadrupole components change sign (red curve). Arrows indicate the typical trajectories of reversals. The time series of the dipole and quadrupole for $\delta q=5.3\%$ is shown in Fig. 5 (b). It is found systematically that during the reversal process, the dipole reverses first, and the quadrupole follows slightly later. A movie of this simulation can be found in the Supplementary Material supp. Reversals follow the opposite trajectory when the amplitude of the heterogeneous heat flux is negative (blue curve). A long evolution of the dipolar angle and the dipolar strength $f_{\mathrm{dip}}$ for $\delta q=5.3\%$ is shown in Fig. 6. In the period of $120$ dipole diffusion times, we observe several non-periodic chaotic reversals of the magnetic field. During the reversal, the dipolar strength decreases from $f_{\mathrm{dip}}\approx 0.9$ to a value $f_{\mathrm{dip}}<0.5$, consistent with the reduction of the dipole component during this event shown in Fig. 5(b). The high value of $f_{\mathrm{dip}}$ between reversals can be attributed to the presence of the SSL, allowing for a strong dipolar magnetic field that reverses, in contrast with the magnetic reversals observed in the absence of a layer in the multipolar regime, dominated by high order modes. We also verified the robustness of these solutions by introducing small variations (up to $10\%$) in both the Rayleigh number and the amplitude of the heterogeneous heat-flux pattern $Y_{3}^{0}$. In these cases, we recover reversals with similar properties to those shown in Figs. 5 and 6.
@@
We can also consider a linear combination of the two patterns $Y_{1}^{0}$ and $Y_{3}^{0}$ as a heterogeneous heat flux at the CMB. Figure 7 shows the evolution of the magnetic field in the D-Q phase space for a linear combination of these two patterns, both with the same amplitude and a total flux variation $\delta q=1.75\%$. A movie of this simulation can be found in the Supplementary Material supp. This heterogeneous heat flux leads to polarity reversals of the magnetic field, with a trajectory that is qualitatively similar to the one described by $Y_{3}^{0}$, but here the time delay between the reversal of the dipole and quadrupole is larger, and the trajectories deviate significantly from the origin. The arrows in the figure show the direction of the reversal. The phenomenology observed here differs from the one described in the absence of a SSL, where both the dipole and quadrupole components fluctuate around the origin, as shown in the inset of Fig. 7(a). We performed also a set of simulations (not shown here) to explore if it is possible to recover reversals in the absence of a SSL and for $\mathrm{Ra}/\mathrm{Ra}_{c}=5$. For high amplitudes of $\delta q$, we do not recover dynamo solutions as a consequence of a high symmetry breaking of the flow. For low amplitudes of $\delta q$ we do observe an enhancement of the quadrupolar component of the magnetic field. In the next section, we provide an interpretation of this phenomenology based on a low-dimensional model and the difference of the dipole and quadrupole growth rates.
@@ 5 Discussion
Figure 7: Evolution of the dipole and quadrupole magnetic field at the CMB for a linear combination of the heterogeneous heat flux patterns $Y_{1}^{0}+Y_{3}^{0}$ with an amplitude $\delta q=1.75\%$. The arrows indicate the orientation of reversals. The inset shows a typical simulation with reversals in a fully convective domain. All simulations correspond to $\mathrm{E}=10^{-3}$, $\Pr=1$, $\mathrm{Pm}=10$, and $\mathrm{Ra}/\mathrm{Ra}_{c}=25$.
The phenomenology presented in the previous section, involving the emergence of hemispheric dynamos and polarity reversals due to an equatorial symmetry breaking, can be interpreted using low-dimensional models. These models consider that the axial dipole $D$ and axial quadrupole $Q$ are the two dominant components of the magnetic field, which interact in a non-linear manner coupled by the hydrodynamic properties of the system. Taking into account the $\bm{B}\rightarrow-\bm{B}$ invariance of the MHD equations, this model is written as (Petrelis2008; Petrelis2009)
@@
where all the parameters $C_{ij}$, $\alpha$, $\beta$ and $\sigma_{D,Q}$ are real numbers. In particular, the coefficients $\sigma_{D}$ and $\sigma_{Q}$ determine the dipole and quadrupole growth rates in the limit in which these two modes are decoupled and the non-linear interactions are negligible. When $\sigma_{D}$ and $\sigma_{Q}$ are similar, an equatorial symmetry breaking will result either in periodic polarity reversals or in hemispheric dynamos (Gallet2009). In the full model, stochastic term is introduced to account for the effects of turbulence (Petrelis2008). Alternatively, Eqs (18) can be coupled to a third mode modelling the asymmetric velocity mode generated by the heterogeneous heat flux, resulting in a deterministic model for chaotic inversions (Gissinger2012a).
@@
It has been shown that such low-dimensional models based on dipole-quadrupole interactions can reproduce fairly well the magnetic field reversals observed on Earth (Petrelis2009; Petrelis2011) and in experimental dynamos (Berhanu2007; Gissinger2010a). However, this behaviour has never been observed, so far, in numerical modelling of the geodynamo, even in the presence of a heterogeneous heat flux (Gissinger2012). Interestingly, it is therefore the presence of an SSL in the present simulations that gives rise to this long-anticipated phenomenology. This could be a consequence of the attenuation of fast modes of the magnetic field due to the skin effect in the SSL. A more appealing interpretation involves the relative growth rates of the dipole and quadrupole components and their non-linear interaction. To understand the behaviour of these quantities, we perform DNS of the kinematic dynamo model, which corresponds to the same system of equations (1)-(4) eliminating the Lorentz force from Eq. (1). We point out that the velocity field still evolves in time under the same boundary conditions (including the heterogeneous CMB heat flux) as in the self-consistent simulations, but now independently of the magnetic field. The magnetic field therefore responds passively to the temporal fluctuations of this flow. We first focus on the hydrodynamic properties of the system, in particular on the equatorial symmetry of the flow as a function of the Rayleigh number. Figure 8(a) shows the relative antisymmetric kinetic energy for different values of the size of the SSL defined in Eq. (17). For low values of $\mathrm{Ra}/\mathrm{Ra}_{c}$, this quantity is zero indicating that the flow is symmetric with respect to the equator. In this regime, the dipole and quadrupole components are independent as they are decoupled. At around $\mathrm{Ra}=10\mathrm{Ra}_{c}$, the antisymmetric kinetic energy undergoes a transition and the equatorial symmetry of the flow is broken. When this happens, $D$ and $Q$ are coupled. When the two modes have similar growth rates, this favours magnetic reversals. We compute the dipole $\sigma_{D}$ and quadrupole $\sigma_{Q}$ growth rates as
$D(t)=D_{0}\exp(\sigma_{D}t),$ (19)
$Q(t)=Q_{0}\exp(\sigma_{Q}t).$ (20)
It is interesting to note that this transition occurs simultaneously with the dipolar-multipolar transition in the fully convective system, and explains why magnetic reversals are obtained in this regime. This result is consistent with Garcia2017, where it was shown that the loss of dipolarity is associated to a breaking of the equatorial symmetry of the flow.
@@
More interestingly, we now focus on the effect of the SSL on the kinematic growth rates of the dipole and quadrupole modes below the symmetry breaking threshold, in the regime where the two linear growth rates remain decoupled. Figure 8(b) shows the dipole and quadrupole growth rates for $\mathrm{Ra}=7\mathrm{Ra}_{c}$ for different values of $\mathcal{H}_{s}$. When the size of the SSL increases, the relative difference between the kinematic growth rates of the dipolar and quadrupolar modes decreases. This behaviour can be understood with the prediction by Proctor1977, who showed that in $\alpha^{2}$-dynamos surrounded by a perfectly conducting, motionless outer layer, a degeneracy arises between the growth rates of the dipole and quadrupole components. In this context, the SSL beneath the core-mantle boundary can be interpreted as an effective conducting boundary, leading to a similar degeneracy as its thickness increases (Favier2013). These conditions strongly favour the emergence of the low-dimensional dynamics described above: as the dipole and quadrupole growth rates $\sigma_{D}$ and $\sigma_{Q}$ approach one another, even a weak symmetry breaking induces a spectacularly strong coupling between the two modes, enhancing the dynamics predicted by low-dimensional models - including both chaotic reversals and the development of hemispherical magnetic fields.
@@ 6 Conclusions
In this work, we studied the effects of a stably-stratified layer (SSL) below the core-mantle boundary (CMB) on the morphology of the magnetic field, with particular focus on polarity reversals. Using direct numerical simulations of the incompressible magnetohydrodynamics equations in a rotating spherical shell driven by convection, we showed that the SSL enhances the dipolar strength of the magnetic field at the CMB due to the skin effect, that attenuates rapidly fluctuating fields. As a result, the transition from dipolar to multipolar dynamos occurs at higher values of the Rayleigh number when the SSL is present, and the transition becomes more abrupt.
@@
By imposing both an SSL and a heterogeneous heat flux at the CMB to break the equatorial symmetry of the flow, we observed a variety of dynamo regimes, including hemispheric dynamos and polarity reversals. Hemispheric dynamos arise from an increase in the quadrupolar component of the magnetic field, which is due to the imposed heterogeneous heat flux and the resulting weak equatorial symmetry breaking of the flow. For the parameters studied in this work, we found that the simplest heat flux pattern, described by the spherical harmonic $Y_{1}^{0}$, controls the hemisphericity of the magnetic field. In contrast, a more complex pattern $Y_{3}^{0}$ is responsible for triggering polarity reversals. When a linear combination of these two patterns is imposed, reversals follow a clear trajectory in the dipole-quadrupole (DQD-Q) phase space, with the dipole reversing systematically before the quadrupole. This reversal behaviour differs from that observed in fully convective systems within the multipolar regime, where both the dipole and quadrupole components fluctuate without following a well-defined trajectory in the DQD-Q plane. We also highlight that between the reversals, the dipolar strength of the magnetic field takes relatively large values $f_{\mathrm{dip}}\approx 0.9$, that are typically not observed in simulations.
@@
We interpret these results and the role of the SSL by means of kinematic dynamo simulations. In fully convective dynamos, we associate the dipolar-multipolar transition with a hydrodynamical instability - controlled by the Rayleigh number - that breaks the equatorial symmetry of the flow. When the flow remains symmetric with respect to the equatorial plane, the dipole and the quadrupole become decoupled, and the growth rates of the dipole $\sigma_{D}$ and the quadrupole $\sigma_{Q}$ approach each other as the thickness of the SSL increases. This behaviour is consistent with the prediction by Proctor1977, who showed that in $\alpha^{2}$-dynamos surrounded by a perfect conductor, a degeneracy arises between the dipolar and quadrupolar growth rates.
The mechanism based on the interaction of the dipole and quadrupole components of the magnetic field provides an explanation for the mechanism of magnetic reversals observed in the geodynamo. It explains the observations of Garcia2017, where it was shown that the loss of dipolarity is associated to a breaking of the equatorial symmetry of the flow. Recent studies of Aubert2025 and Jones2025 provide other alternative for which the geodynamo model can reproduce Earth-like magnetic reversals. In particular, Aubert2025 provide a kinematic mechanism controlled by the relative strength of subsurface upwelling and horizontal circulation at the core surface. In Jones2025, the authors find a regime of parameters in which the magnetic field reverses even in the low inertia limit by fixing the ratio between $\mathrm{Pm}$ and $\Pr$. Our results propose a different mechanism based on a low-dimensional model used to explain reversals in dynamo experiments Petrelis2009; Petrelis2008. The simulations showing reversing solutions in this work operate close to the dipolar-multipolar transition, but we expect this mechanism to apply also for low-inertia systems.
@@
This interpretation of the SSL, and its critical role in reducing the difference between dipolar and quadrupolar growth rates ($\sigma_{D}-\sigma_{Q}$), provides an elegant resolution to a long-standing question regarding the robustness of magnetic field reversals. Indeed, despite considerable variability in core conditions - such as inner-core growth, fluctuations in the heat flux, and other long-term evolutionary processes - the presence of a substantial SSL acts as a stabilizing factor. By maintaining the near-degeneracy of $\sigma_{D}$ and $\sigma_{Q}$, it ensures that the low-dimensional dynamics driven by the dipole-quadrupole interaction remains robust, even in the presence of other complex small scale modes. Consequently, even small symmetry breaking effects become sufficient to sustain the observed reversal behaviour over geological timescales, making Earth’s dynamo remarkably resilient to changes in its internal properties. Finally, we remark the results shown in this work were obtained for a relatively high value of the Ekman number of $\mathrm{E}=10^{-3}$ and the magnetic Prandtl number $\mathrm{Pm}=10$, which are quite far from realistic values for the Earth of $\mathrm{E}=10^{-15}$ and $\mathrm{Pm}=10^{-6}$. Further studies could investigate the dependence of this phenomenology with the Ekman number and also the magnetic Prandtl number.